### v1

# Laplacian Spectrum and Domination in Trees

Abstract
For a finite simple undirected graph $G$, let $\gamma(G)$ denote the size of a smallest dominating set of $G$ and $\mu(G)$ denote the number of eigenvalues of the Laplacian matrix of $G$ in the interval $[0,1)$, counting multiplicities. Hedetniemi, Jacobs and Trevisan [Eur. J. Comb. 2016] showed that for any graph $G$, $\mu(G)\leqslant\gamma(G)$. Cardoso, Jacobs and Trevisan [Graphs Combin. 2017] asks whether the ratio $\gamma(T)/\mu(T)$ is bounded by a constant for all trees $T$. We answer this question by showing that this ratio is less than $4/3$ for every tree. We establish the optimality of this bound by constructing an infinite family of trees where this ratio approaches $4/3$. We also improve this upper bound for trees in which all the vertices other than leaves and their parents have degree at least $k$, for every $k\geqslant 3$. We show that, for such trees $T$, $\gamma(T)/\mu(T)<1+1/((k-2)(k+1))$.
Keywords.
Laplacian Spectrum, Domination Number, Trees.
Mathematics Subject Classification.
05C50, 05C69, 15A18, 15A42.
1 Introduction
Let $G$ be a simple undirected graph on the vertex-set $V(G)=\{v_{1},\ldots,v_{n}\}$. The adjacency matrix $A(G)=[a_{ij}]$ of $G$ is an $n\times n$ matrix where the entry $a_{ij}$ is $1$ if there is an edge between vertices $v_{i}$ and $v_{j}$, and 0 otherwise. The Laplacian matrix of $G$ is L(G)=D(G)A(G)L(G)=D(G)-A(G), where $D(G)$ is a diagonal matrix whose $i$-th diagonal entry is the degree of $v_{i}$. The multi-set of eigenvalues of $L(G)$ is termed the Laplacian spectrum of $G$. The Laplacian spectrum of $G$ is always a subset of $[0,n]$. Let $m_{G}I$ denote the number of Laplacian eigenvalues of $G$ within a real interval $I$, counting multiplicities. The value of $m_{G}I$ for different choices of $I$ bound many combinatorial parameters of the graph $G$. In this paper we study how $m_{G}[0,1)$ bounds the domination number of trees. Henceforth, we abbreviate $m_{G}[0,1)$ as $\mu(G)$.
A set of vertices $D$ in $G$ is called dominating if every vertex of $G$ is either in $D$ or is adjacent to a vertex in $D$. The domination number $\gamma(G)$ is the size of a smallest dominating set in $G$. The connection between Laplacian spectrum and the domination number of a graph was first investigated by Hedetniemi, Jacobs, and Trevisan hedetniemi2016domination in 2016 who established the following lemma.
Lemma 1 (hedetniemi2016domination ).
For every graph $G$, $\mu(G)\leqslant\gamma(G)$.
This study was continued by Cardoso, Jacobs, and Trevisan cardoso2017laplacian . Along with $\mu(G)=m_{G}[0,1)$, they also studied the dependence of $\gamma(G)$ on $\nu(G)=m_{G}[2,n]$. They proved that $\nu(G)\geqslant\gamma(G)$ for graphs without isolated vertices. They further demonstrated that ratios $\gamma(G)/\mu(G)$ and $\nu(G)/\gamma(G)$ can become arbitrarily large. In contrast, they showed that for trees and graphs which can be made acyclic by removing a bounded number of edges, the ratio $\nu(G)/\gamma(G)$ is bounded. The study concluded with the problem of determining whether the ratio $\gamma(T)/\mu(T)$ is bounded for trees. The main contribution of this paper is the following affirmative answer to this question.
Theorem 2.
For every tree $T$,
$1\leqslant\frac{\gamma(T)}{\mu(T)}<\frac{4}{3}.$
We show the optimality of this upper bound by constructing an infinite family of trees for which the ratio approaches $4/3$ (c. f. Section 4.) So far, we did not know of any family of trees where this ratio was larger than $25/24$ cardoso2017laplacian . We also show that the bound can be improved for trees in which all deep vertices (Definition 1) have high degree.
Definition 1.
A vertex of a tree is called a leaf if its degree is $1$, penultimate if it is adjacent to a leaf and deep if it is neither. We denote the number of penultimate vertices of a tree $T$ by $p(T)$.
Theorem 3.
If $T$ is a tree in which all deep vertices have a degree at least $k$, $k\geqslant 3$,
$\gamma(T)<\left(1+\frac{1}{(k-2)(k+1)}\right)p(T).$
Since $p(T)\leqslant\mu(T)$ (Lemma 7), this result improves the upper bound in Theorem 2 for every $k\geqslant 3$. The upper bound in Theorem 2 is established by analysing an extended version (Algorithm 2) of an algorithm (Algorithm 1) by Braga, Rodrigues, and Trevisan braga2013distribution . While Algorithm 1 computes the number of Laplacian eigenvalues less than, equal to and greater than a certain value, our extension also constructs a dominating set of the tree enroute. We tried to apply the same technique to improve the upper bound of $2$ on the ratio $\nu(T)/\gamma(T)$ for trees cardoso2017laplacian . However, we discover a family of caterpillars (trees with a dominating path) for which this ratio approaches $2$ (c. f. Section 6).
Background.
We now list some of the known connections between $m_{G}I$ for different choices of $I$ and combinatorial parameters of graphs. In 1990, Grone, Merris, and Sunder grone1990laplacian proved that $m_{G}[0,1)\geqslant q(G)$, where $q(G)$ is the number of vertices in $G$ which are not pendant but adjacent to a pendant vertex. In 2013, Braga, Rodrigues, and Trevisan braga2013distribution demonstrated that for any tree $T$ with more than one vertex, the inequality $m_{T}[0,2)\geqslant\lceil\frac{n}{2}\rceil$ holds. They further conjectured that $m_{T}[0,\overline{d})\geqslant\lceil\frac{n}{2}\rceil$, where $\overline{d}$ represents the average degree of $T$. This conjecture was subsequently confirmed by Jacobs, Oliveira, and Trevisan jacobs2021most , and Sin sin2020number independently in 2021. In 2022, Ahanjideh et al. ahanjideh2022laplacian established that, for a connected graph $G$, $m_{G}[0,1)\leqslant\alpha(G)\leqslant m_{G}[0,n-\alpha(G)]$, where $\alpha(G)$ is the independence number of $G$. They also characterized graphs for which $m_{G}[0,1)=\alpha(G)=\frac{n}{2}$. Additionally, they discovered that for connected triangle-free or quadrangle-free graphs $G$, $m_{G}(n-1,n]\leqslant 1$, and characterized graphs such that $m_{G}[0,1]=n-1$ or n2n-2. Further, they showed that in connected graphs, $m_{G}(n-1,n]\leqslant\chi(G)-1$, where $\chi(G)$ is the chromatic number of $G$. Later, Akbari et al. akbari2023classification , and Cui and Ma cui2024laplacian derived several bounds on $m_{G}I$ based on various graph parameters. Also in 2024, Xu and Zhou explored the Laplacian eigenvalue distribution relative to the diameter of the graph xu2024diameter ; xu2024laplacian . Recently, Zhen, Wong, and Songnian Xu zhen2025laplacian and Leyou Xu and Zhou xu2025girth studied how the Laplacian eigenvalue distribution relates to the girth of a graph, while Akbari, Alaeiyan, and Darougheh akbari2025laplacian studied it in terms of degree sequence. In 2025, Guo, Xue and Liu guo2025laplacian proved that for a tree $T$ with diameter $d$, $m_{T}[0,1)\geqslant\frac{d+1}{3}$, and characterized trees that achieve this bound. Moon and Park moon2025laplacian gave a lower bound on $m_{G}[0,1)$ for unicyclic graphs in terms of its diameter and girth.
Organization of the paper.
Section 2 examines an algorithm by Braga, Rodrigues, and Trevisan braga2013distribution which can be used to count the number of Laplacian eigenvalues of a tree in any given interval. This is an extension of an algorithm by Jacobs and Trevisan jacobs2011locating , which does the same for the adjacency spectrum of trees. Section 3 details the proof of Theorem 2 and Section 4 demonstrates the tightness of the same by constructing an infinite family of trees where the ratio $\gamma(T)/\mu(T)$ approaches $4/3$. Section 5 proves Theorem 3. Reading the proof of Theorem 3 first may help in following the proof of Theorem 2 more easily. Section 6, presents an infinite family of trees for which the ratio $\nu(T)/\gamma(T)$ approaches $2$. Section 7 presents some final remarks and discusses a future direction.
2 Sylvester’s Law of Inertia
The inertia of a matrix is the number of its positive, negative, and zero eigenvalues, counting multiplicities. Two matrices $A$ and $B$ over a field are called congruent if there exists an invertible matrix $P$ over the same field such that $P^{T}AP=B$.
Lemma 4 (Sylvester’s Law of Inertia horn2012matrix ).
Two real symmetric matrices are congruent if and only if they have the same inertia.
Sylvester’s Law of Inertia helps one to count the number of eigenvalues of a real symmetric matrix $M$ in any interval $I$. For example, if $I=[a,b)$, the number of eigenvalues of $M$ can be found from the inertia of the matrices MaIM-aI and MbIM-bI. In particular $\mu(G)$ is equal to the number of negative eigenvalues of L(G)IL(G)-I. We use Algorithm 1 by Braga, Rodrigues, and Trevisan braga2013distribution which determines the inertia of the matrix $L(T)+\alpha I$ for a tree $T$ by doing a single bottom-up scan of $T$ without using any matrix computations. This algorithm outputs a diagonal matrix that is congruent to $L(T)+\alpha I$. Each entry in the diagonal of the matrix corresponds to the values of the respective vertices at the end of the algorithm.
Algorithm 1 braga2013distribution
Input: Rooted Tree $T$, scalar $\alpha$
Output: Diagonal matrix $D$ congruent to $L(T)+\alpha I$
1:Initialize $f(v):=d(v)+\alpha$ for all vertices $v$
2:Order vertices as $v_{1},\dots,v_{n}$ from the lowest level to the highest level (root)
3:for $k=1$ to $n$ do
4:  if $v_{k}$ is a leaf then
5:   continue
6:  else if $f(c)\neq 0$ for all children $c$ of $v_{k}$ then
7:   $f(v_{k}):=f(v_{k})-\sum\limits_{\begin{subarray}{c}c\end{subarray}}\frac{1}{f(c)}$, where the sum over all children of $v_{k}$.
8:  else
9:   select one child $v_{j}$ of $v_{k}$ for which $f(v_{j})=0$
10:   $f(v_{k}):=-\frac{1}{2}$, $f(v_{j}):=2$
11:   if $v_{k}$ has a parent $v_{t}$ then
12:     remove the edge $v_{k}v_{t}$
13:   end if
14:  end if
15:end for
16:Let $D$ be a diagonal matrix with $d_{ii}=f(v_{i})$ for each vertex $v_{i}$ in $T$
17:return $D$
Theorem 5 (jacobs2011locating ).
Let $D$ be the diagonal matrix produced by Algorithm 1 applied to a tree $T$ rooted at any of its vertices and scalar $\alpha=-\beta$. Then, $m_{T}(\beta,n]$, $m_{T}[0,\beta)$$m_{T}[\beta]$ are equal to the number of positive, negative, and zero values on the diagonal of $D$.
Definition 2.
The final value $f(v)$ assigned to a vertex $v$ by Algorithm 1 is referred to as the inertia of $v$. The vertex $v$ is called negative if its inertia $f(v)$ is negative. We will say that a vertex $u$ is a zero-child of its parent $v$ if $f(u)=0$ when Algorithm 1 starts to process $v$, irrespective of whether the inertia of $u$ is 0 or $2$.
The main tool in the paper is an analysis of an extension of this algorithm (Algorithm 2). We will make repeated use of the following direct corollary of Theorem 5 since our extended algorithm simulates a run of Algorithm 1 with $\alpha=-1$.
Corollary 6.
For a tree $T$, $\mu(T)$ is equal to the number of negative vertices in $T$ after a run of Algorithm 1 on $T$ with $\alpha=-1$.
3 Proof of the Main Result
In 1990, Grone et al. grone1990laplacian proved that for any graph $G$, $\mu(G)\geqslant q(G)$ where $q(G)$ is the number of vertices in $G$ which are not pendant but adjacent to a pendant vertex. The following Lemma follows from the same. We also give a direct proof for trees using Algorithm 1 and Theorem 5.
Lemma 7.
For any tree $T,\ \mu(T)\geqslant p(T)$.
Proof.
When $\alpha$ is 1-1, Algorithm 1 initializes all the leaf vertices with a value of 0 and hence every penultimate vertex gets an inertia of 1/2-1/2. ∎
3.1 Handling Adjacent Two-degree Vertices
In this section, we show that any upper bound (of at least $1$) on $\gamma(T)/\mu(T)$ that one can prove on trees with no adjacent $2$-degree vertices can be extended to all trees. This helps us in simplifying our later proofs.
Definition 3.
A path in a graph $G$ is called clean if all its internal vertices have degree exactly two in $G$.
Definition 4.
The contraction of a set of vertices $S\subseteq V(G)$ in $G$ replaces the vertices in $S$ with a single new vertex, which is adjacent to the neighbors of all the vertices in $S$.
Lemma 8.
Let $T^{\prime}$ be a tree obtained from a tree $T$ by contracting a clean path of length three. Then
$\frac{\gamma(T)}{\mu(T)}\leqslant\frac{\gamma(T^{\prime})}{\mu(T^{\prime})}.$
Proof.
We shall first show that $\mu(T)=\mu(T^{\prime})+1$. To analyze Algorithm 1 without splitting cases when a vertex gets an $f$-value of zero, we will set $1/0=\infty$ and work with the extended real number system. This only results in the following changes in Algorithm 1. If a vertex $v$ has a zero-child, then instead of setting f(v)=1/2f(v)=-1/2, we will get $f(v)=$-$\infty$. The modified algorithm will also change the inertia of one of the children from zero to $2$. However, we need not remove the edge from $v$ to its parent, since the parent gets a contribution of $1/f(v)=-1/\infty=0$ from $v$. This simulates the effect of deleting the edge from $v$ to its parent. Therefore, this alternate analysis doesn’t affect the final sign of $f(v)$ for any vertex $v$. This approach is used only in the proof of this Lemma.
Let $(a,b,c,d)$ be a clean path in $T$ with $b$ being a child of $a$. Let $b_{1},\dots,b_{l}$ be the other children of $a$, and $e_{1},\dots,e_{k}$ be the children of $d$. After contracting $(a,b,c,d)$ into a single vertex $e$, the children of $e$ are $b_{1},\dots,b_{l},e_{1},\dots,e_{k}$. We will use $f_{T}$ and $f_{T^{\prime}}$ respectively to denote the inertia computed on $T$ and $T^{\prime}$. Since $f_{T}(b)=1-\frac{1}{f_{T}(c)}$ and $f_{T}(c)=1-\frac{1}{f_{T}(d)}$, we see that $1-\frac{1}{f_{T}(b)}=f_{T}(d)$. Using this (in the third step) we can see that
$f_{T}(a)$ $=(l+1)-\left(\frac{1}{f_{T}(b)}+\sum_{j=1}^{l}\frac{1}{f_{T}(b_{j})}\right)$
$=\left(1-\frac{1}{f_{T}(b)}\right)+\left(l-\sum_{j=1}^{l}\frac{1}{f_{T}(b_{j})}\right)$
$=f_{T}(d)+\left(l-\sum_{j=1}^{l}\frac{1}{f_{T}(b_{j})}\right)$
$=\left(k-\sum_{i=1}^{k}\frac{1}{f_{T}(e_{i})}\right)+\left(l-\sum_{j=1}^{l}\frac{1}{f_{T}(b_{j})}\right)$
$=(k+l)-\left(\sum_{i=1}^{k}\frac{1}{f_{T}(e_{i})}+\sum_{j=1}^{l}\frac{1}{f_{T}(b_{j})}\right)$
$=f_{T^{\prime}}(e).$
Thus, this contraction does not affect the execution of the algorithm in other parts of the tree. Hence, all vertices of $T^{\prime}$ other than $e$ retain their inertia as in $T$. One can verify that either $f(d)<0$ or $f(c)<0$ (when $0\leqslant f(d)<1$) or $f(b)<0$ (when $f(d)\geqslant 1$). Therefore, $\mu(T)=\mu(T^{\prime})+1$.
Next we show that $\gamma(T)\leqslant\gamma(T^{\prime})+1$. Let $D^{\prime}$ be a minimum dominating set of $T^{\prime}$. If $e\in D^{\prime}$, then $D=(D^{\prime}\setminus\{e\})\cup\{a,d\}$ is a dominating set of $T$. If $e\not\in D^{\prime}$ then $D=D^{\prime}\cup\{x\}$, where $x$ is the vertex in $\{b,c\}$ which is at a distance of $3$ from the vertex in $D^{\prime}$ which dominates $e$ is a dominating set of $T$. Therefore, $\gamma(T)\leqslant\gamma(T^{\prime})+1$.
Hence
$\frac{\gamma(T)}{\mu(T)}\leqslant\frac{\gamma(T^{\prime})+1}{\mu(T^{\prime})+1}\leqslant\frac{\gamma(T^{\prime})}{\mu(T^{\prime})},$
where the second inequality is a simple observation from the fact that $\gamma(T^{\prime})/\mu(T^{\prime})\geqslant 1$ (Lemma 1). ∎
3.2 A New Dominating Set Algorithm
Let $\mathcal{T}$ denote the set of all finite trees with no adjacent $2$-degree vertices. We design a new algorithm (Algorithm 2) which simulates a run of Algorithm 1 with $\alpha=-1$ on a tree $T\in\mathcal{T}$ and constructs a subset $D$ of $V(T)$ in the same pass using a weight propagation. We will then show that $D$ is a dominating set of size at most $\mu(T)+\frac{1}{3}(p(T)-1)$. Note that, since Algorithm 2 simulates Algorithm 1 with $\alpha=-1$, the number of negative vertices in $T$ at the end of Algorithm 2 is equal to $\mu(T)$. Recall that, by Lemma 8, it is enough to establish an upper bound for trees in $\mathcal{T}$ in order to obtain an upper bound for all trees.
Throughout this section, we fix an arbitrary tree $T$ from $\mathcal{T}$ and the set computed by Algorithm 2 on $T$ as $D$. We start by making two observations on the run of Algorithm 2 on $T$ before proving the key result that $D$ is a dominating set of $T$ (Lemma 11).
Algorithm 2
Input: Tree $T\in\mathcal{T}$
Output: A dominating set of $T$.
1:Root $T$ at a penultimate vertex.
2:Initialize f(v):=d(v)1f(v):=d(v)-1 for all vertices $v$, $D=\emptyset$ and
$w(v)=\begin{cases}4/3,&\text{if $v$ is a penultimate vertex}\\ 0,&\text{otherwise}\end{cases}$
3:Process the non-leaf vertices bottom-up so that each vertex is processed after all its children (all leaves are assumed to be processed).
4:for a non-leaf vertex $v$ do
5:  if $f(c)\neq 0$ for all children $c$ of $v$ then
6:   $f(v):=f(v)-\sum\limits_{\begin{subarray}{c}c\end{subarray}}\frac{1}{f(c)}$, where the sum over all children of $v$.
7:  else
8:   select one child $c$ of $v$ for which $f(c)=0$
9:   $f(v):=-\frac{1}{2}$, $f(c):=2$
10:   $w(v)=w(v)+1$
11:   if $v$ has a parent $u$, remove the edge vuvu
12:  end if
13:  if [($w(v)\geqslant 4/3$) or ($f(v)=0$ and $w(v)\geqslant 1$)] and at least one child of $v$ is not dominated by $D$ then
14:   $D=D\cup\{v\}$ and push a weight of (w(v)1w(v)-1) to $v$’s parent.
15:  else
16:   Push a weight of $w(v)$ to $v$’s parent
17:  end if
18:end for
19:return $D$
Observation 9.
Every non-leaf vertex $v$ pushes a weight of at least $1/3$ to its parent unless $v$ is added to $D$ and $v$ is a zero-child of its parent.
Proof.
We prove this by induction on the distance $d_{v}$ of $v$ to the nearest leaf below it. The base case is when $v$ is a penultimate vertex ($d_{v}=1)$. In this case, $v$ pushes a weight of 4/31=1/34/3-1=1/3 to its parent, confirming the claim. Let $v$ be a vertex with $d_{v}>1$ and let us assume that for every vertex $u$ with $d_{u}<d_{v}$, the claim is true. In particular, the claim is true for every child of $v$.
If every child of $v$ is a zero-child and all of them are added to $D$, then f(v)=1/2f(v)=-1/2 and $v$ is not added to $D$. Hence $w(v)\geqslant 1$ is pushed to its parent. Otherwise, $v$ gets a total weight of at least $1/3$ from its child(ren) which is pushed to its parent unless $v$ is added to $D$. The vertex $v$ will be added to $D$ only if either $w(v)\geqslant 4/3$, in which case a weight of at least $1/3$ is pushed to its parent or if $f(v)=0$ when processing $v$, in which case $v$ is exempted from pushing any weight to its parent $u$ as it is a zero-child of $u$ which is added to $D$. ∎
Observation 10.
If $v$ is a negative vertex with no zero-child, then it pushes a weight of at least $2/3$ to its parent.
Proof.
Since $v$ is a negative vertex, $w(v)\geqslant 1$. If $v$ has at least two children, then according to Observation 9, it receives a weight of at least $1/3$ from each child since none of them is a zero-child. Consequently, the total weight of $v$ is at least $5/3$, allowing it to push a weight of at least $2/3$ to its parent.
Suppose $v$ has only one child, say $x$, then $x$ is not a leaf since $v$ has no zero-child. Moreover, $x$ must have at least two children since $T$ does not have two adjacent $2$-degree vertices. If $x$ is dominated when $v$ is being processed, then $v$ pushes all its weight to its parent. Otherwise, no child of $x$ is in $D$, and so all of them push at least $1/3$ weight to $x$, which in turn pushes this weight to $v$. As a result, $v$ has a weight of at least $5/3$ and thus pushes at least $2/3$ to its parent. ∎
Lemma 11.
$D$ dominates $T$.
Proof.
For a vertex $v\in V(G)$, the notation $N[v]$ denotes the set of vertices adjacent to $v$, including $v$ itself. Since all the penultimate vertices are added to $D$, all the penultimates (which includes the root of $T$), and the leaves are dominated. Let $v$ be any non-root deep vertex of $T$ and $u$ be its parent. We will verify that the total weight entering $N[v]$ (including both initialization and propagation) is at least $4/3$, or at least $1$ with $f(u)=0$ when $u$ is being processed. This ensures the inclusion of at least one vertex of $N[v]$ in $D$, guaranteeing that $v$ is dominated by $D$.
We call the children of any vertex in $N[v]$ that do not belong to $N[v]$ as feeders of $v$. These feeders consist of the siblings and grandchildren of $v$. The parents of these feeders will be called feeder inputs of $v$. Every deep vertex in $T$ has either at least two children or a sibling. Hence, $v$ has at least two feeder inputs and two feeders. If any of the feeders is a zero-child, then its parent is negative and gets an initial weight of $1$. The remaining feeder inputs (at least one exists) gets a weight of at least $1/3$ from its feeders (Observation 9). Hence, $N[v]$ gets a total weight of at least $4/3$ and we are done. So we can assume henceforth that no feeder is a zero-child. Hence, every feeder pushes a weight of at least $1/3$ to its parent (Observation 9). Let $F$ denote the set of feeders of $v$. We split the analysis based on $\lvert F\rvert$.
Case 1.
$\lvert F\rvert\geqslant 4$
Since each feeder pushes a weight of at least $1/3$ to its parent, $N[v]$ gets a total weight of at least $4/3$.
Case 2.
$\lvert F\rvert=3$
If $v$ has a feeder $z$ which is negative and has no zero-child, then $z$ pushes a weight $2/3$ to its parent (Observation 10). Hence $N[v]$ gets a total weight of at least $4/3$ and we are done. So we can assume that every feeder of $v$ either has a positive inertia or has a zero-child. We call a feeder vertex which either has a positive inertia or a zero-child as pseudo-positive. Since $\lvert F\rvert=3$, the feeders push a weight of at least $1$ to $N[v]$. It is enough to show that there is either a negative vertex in $N[v]$ or $f(u)=0$ when $u$ is processed. The key observation is that if all the children of a vertex $x$ are pseudo-positive, then $f(x)\leqslant deg(x)-1$. For every child $x$ of $v$, since all the children of $x$ are pseudo-positive (being feeders), we conclude that $f(x)\in(0,deg(x)-1]$.
There are four subcases to consider - two when $v$ has one sibling and two grandchildren, and two when $v$ has no siblings but has three grandchildren (See Figure 1).
  1. 1.
    $v$ has one sibling and two children $x_{1}$, $x_{2}$ with one child each.
    In this case $f(x_{1}),f(x_{2})\in(0,1]$ and hence $f(v)\leqslant 0$. Thus either $v$ or $u$ is negative.
  2. 2.
    $v$ has one sibling $v_{1}$ and one child $x_{1}$ with two children.
    In this case $f(x_{1})\in(0,2]$ and $f(v)\leqslant\frac{1}{2}$. If $f(v)\leqslant 0$, then we are done. Otherwise, we will use the fact that $v_{1}$ is pseudo-positive. If $v_{1}$ has a zero-child, then $f(u)=2-\frac{1}{f(v)}\leqslant 0$.
    If $f(v_{1})>0$, $f(u)=2-\frac{1}{f(v)}-\frac{1}{f(v_{1})}<0$.
  3. 3.
    $v$ has no siblings but three children $x_{1}$, $x_{2}$, $x_{3}$ with one child each.
    In this case $f(x_{i})\in(0,1],\forall i\in\{1,2,3\}$ and hence $f(v)\leqslant 0$.
  4. 4.
    $v$ has no siblings but two children $x_{1}$, $x_{2}$, with one child for $x_{1}$ and two for $x_{2}$.
    In this case $f(x_{1})\in(0,1]$ and $f(x_{2})\in(0,2]$. Hence $f(v)\leqslant\frac{1}{2}$. So either $f(v)\leqslant 0$ or $f(u)<0$.
[…]
1
[…]
2
[…]
3
[…]
4
Figure 1: Four subcases when $|F|=3$ and all feeders of $v$ are pseudo-positive in the proof of Lemma 11
Case 3.
$\lvert F\rvert=2$
Since $v$ has at least two grandchildren, $v$ has no siblings in this case. Let $x_{1}$, $x_{2}$ be the children of $v$, each with exactly one child. If both grandchildren are negative vertices with no zero-child, then each of them pushes a weight of at least $2/3$ to $N[v]$, and we are done. Hence, at least one of them, say the child of $x_{1}$ is pseudo-positive. In this case, $f(x_{1})\leqslant 1$. If $f(x_{1})\leqslant 0$ or $f(x_{2})\leqslant 0$, we are done and hence we can assume $f(x_{1})\in(0,1]$ and $f(x_{2})>0$. Then, $f(v)<1$. Hence either $f(v)\leqslant 0$ or $f(u)<0$.
From Cases 1, 2, and 3, it follows that the weight entering $N[v]$ is at least $4/3$, or at least $1$ with $f(u)=0$ when $u$ is processed. Thus, at least one vertex from $N[v]$ is included in $D$, which implies that $D$ dominates $v$. ∎
Remark.
As an aside, we note that $D$ is infact a minimum dominating set of $T$. This can be seen by comparing Algorithm 2 with the following greedy algorithm for finding a minimum dominating set in any tree $T$. Begin by setting $D=\emptyset$ and process the vertices of $T$ bottom up. If a vertex $v$ has at least one child that is not dominated by $D$, add $v$ to $D$. A version of this algorithm is analyzed in stackoverflow to show that the process results in a minimum dominating set.
Lemma 12.
$|D|\leqslant\mu(T)+\frac{1}{3}(p(T)-1).$
Proof.
According to Lemma 11, the set $D$ is a dominating set. It is easy to see that and $|D|$ is at most the total initial weight. The root vertex $v$ being a penultimate vertex, gets an initial weight of $4/3$ and retains a weight of at least $1/3$. Hence,
$\lvert D\rvert\leqslant w(T)-\frac{1}{3}=\frac{4}{3}p(T)+(\mu(T)-p(T))-\frac{1}{3}=\mu(T)+\frac{1}{3}(p(T)-1).$
3.3 Completing the Proof of Theorem 2
Let $T\in\mathcal{T}$. We run Algorithm 2 on $T$ to obtain a set $D$. By Lemma 11, $D$ is a dominating set and by Lemma 12 the size of $D$ is bounded above by $\mu(T)+\frac{1}{3}p(T)$. Hence we have
$\gamma(T)<\mu(T)+\frac{1}{3}p(T)\leqslant\mu(T)+\frac{1}{3}\mu(T)=\frac{4}{3}\mu(T),$
where the second inequality follows from Lemma 7.
Thus $\gamma(T)/\mu(T)<4/3$ for any tree in $\mathcal{T}$. From Lemma 8, we can extend this upper bound to all trees. The lower bound for the ratio $\gamma(T)/\mu(T)$ is immediate from Lemma 1. Therefore, for all trees $T$,
$1\leqslant\frac{\gamma(T)}{\mu(T)}<\frac{4}{3}.$
This completes the proof of Theorem 2.
4 Tight Example
Hitherto, the tree with the largest known value of $\gamma/\mu$ was a tree on 65 vertices and it had $\gamma/\mu=25/24$ cardoso2017laplacian . In Figure 3, we show a tree on $11$ vertices with $\gamma=5$ and $\mu=4$. By adding more leaf vertices to the penultimates, we can find an infinite family of trees with $\gamma=5$ and $\mu=4$. Now, we construct an infinite family of trees to show that our upper bound of $4/3$ is tight.
Consider the tree $T_{k}$ illustrated in Figure 3. $T_{k}$, $k\geqslant 2$, has $k$ isomorphic subtrees hanging from the root. We only illustrate the leftmost subtree in the figure. By analyzing the run of Algorithm 1 on $T_{k}$, we can see that the negative vertices are the root vertex $r$ and the penultimate vertices. Hence $\mu(T_{k})=3k+1$. Furthermore, a minimum dominating set of $T_{k}$ must include at least four vertices from each subtree hanging from $r$. Hence $\gamma(T_{k})\geqslant 4k$. It is straightforward to construct a dominating set by picking four vertices from each subtree.
02212-$\frac${1}{2}2212-$\frac$$2-\frac{1}{2}2212$-$\frac${1}{2}22
Figure 2: 11 vertex tree with $\gamma=5$ and $\mu=4$. Inertia of each vertex is indicated.
[…]
Figure 3: Tree with inertia of vertices. The penultimate vertices are indicated by dark circles.
Hence $\gamma(T_{k})=4k$ and
$\frac{\gamma(T_{k})}{\mu(T_{k})}=\frac{4k}{3k+1}.$
As $k$ increases, this ratio approaches $4/3$ showing that the upper bound in Theorem 2 is optimal.
5 Improving the Upper Bound for High-Degree Trees
Let $\mathcal{T}_{k}$, $k\geqslant 3$, denote the set of all finite trees in which every deep vertex has degree at least $k$. We first establish an upper bound for the domination number in terms of the number of its penultimate vertices, and then see that this improves the upper bound of Theorem 2 for such trees.
Consider any tree $T\in\mathcal{T}_{k}$. We run Algorithm 3 on $T$ to obtain a set $D$ and proceed to analyse $D$ as in Section 3.2. This algorithm is much simpler than Algorithm 2 since it does not rely on Algorithm 1 and only the penultimate vertices are initialised with a non-zero weight.
Algorithm 3
Input: $k\geqslant 3$ and $T\in\mathcal{T}_{k}$
Output: A dominating set of $T$.
1:Root $T$ at a penultimate vertex
2:Initialize $D=\emptyset$ and $\epsilon=\frac{1}{(k-2)(k+1)}$,
$w(v)=\begin{cases}1+\epsilon,&\text{if $v$ is a penultimate vertex}\\ 0,&\text{otherwise}\end{cases}$
3:Process the vertices bottom up.
4:for a vertex $v$ do
5:  if $w(v)\geqslant 1+\epsilon$ then
6:   Add $v$ to $D$ and push a weight of (w(v)1w(v)-1) to $v$’s parent
7:  else
8:   Push a weight of $w(v)$ to $v$’s parent
9:  end if
10:end for
11:return $D$
Observation 13.
Every non-leaf vertex pushes a weight of at least $\epsilon$ to its parent.
Proof.
We prove this by induction on the distance $d_{v}$ of $v$ to the nearest leaf below it. The base case is when $v$ is a penultimate vertex ($d_{v}=1)$, in which case, $v$ pushes $\epsilon$ to its parent. Let $v$ be a vertex with $d_{v}>1$ and let us assume that for every vertex $u$ with $d_{u}<d_{v}$, the claim is true. If $v\in D$, then $w(v)\geqslant 1+\epsilon$, ensuring that a weight of at least $\epsilon$ is pushed to its parent. Otherwise, by induction hypothesis, it receives a weight of at least $\epsilon$ from its children (none of which are leaves), and all of it is passed on to its parent. ∎
Lemma 14.
$D$ dominates $T$.
Proof.
Let $v$ be an arbitrary vertex in $T$. We will show that $v$ is dominated by $D$. Since all the penultimate vertices of $T$ are added to $D$, we can assume that no child or grandchild of $v$ is a leaf. We denote parent of $v$ as $u$. If $v$ or any of its children is in $D$, then we are done. If no child of $v$ is in $D$, by Observation 13, each grandchild of $v$ pushes a minimum weight of $\epsilon$ to its parent. Consequently, each child of $v$ gets and pushes to $v$ a weight of at least $(k-1)\epsilon$. Thus $v$ gets a weight of at least $(k-1)^{2}\epsilon$. Since $v$ is not in $D$, this weight is passed to its parent $u$. If $u$ is a penultimate vertex, then we are done. Otherwise, all siblings of $v$ push a minimum weight of $\epsilon$ to $u$ and since there are at least (k2)(k-2) siblings for $v$, we have $w(u)\geqslant(k-1)^{2}\epsilon+(k-2)\epsilon\geqslant 1+\epsilon$, ensuring the inclusion of $u$ in $D$. Therefore $v$ is dominated. ∎
Proof of Theorem 3.
Let $T\in\mathcal{T}_{k}$, $k\geqslant 3$. As the root vertex of $T$ is a penultimate vertex, it retains a weight of at least $\epsilon=1/((k-2)(k+1))$, and from Lemma 14,
$\gamma(T)\leqslant\lvert D\rvert\leqslant w(T)-\epsilon<w(T)=\left(1+\frac{1}{(k-2)(k+1)}\right)p(T).$
Remark.
We can run Algorithm 3 on any tree $T\in\mathcal{T}$ (i.e., trees without adjacent $2$-degree vertices) with $\epsilon=1$. Since every vertex $v$ in $T$ not dominated by a penultimate vertex will have at least two non-leaf grandchildren, either $v$ or one of its child will be in $D$ and hence $D$ will dominate $T$. This, together with Lemma 8, shows that $\gamma(T)/\mu(T)<2$ for every tree $T$. However we needed a much more involved argument to establish the optimal upper bound of $4/3$ (Theorem 2).
Corollary 15.
For any tree $T$ in which all deep vertices have degree at least $k$, $k\geqslant 3$,
$1\leqslant\frac{\gamma(T)}{\mu(T)}<1+\frac{1}{(k-2)(k+1)}.$
Proof.
Follows from Lemma 1, Theorem 3, and Lemma 7. ∎
Remark.
Using Lemma 8, we can extend Corollary 15 to include trees which can be obtained from a tree $T\in\mathcal{T}_{k}$ by replacing one or more edges of $T$ with paths of length 0 or 11 modulo 33.
6 Tightness of the bound $\nu(T)/\gamma(T)<2$ for trees
Recall that $\nu(G)=m_{G}[2,n]$. In this section, we show that the bound $\nu(T)<2\gamma(T)$ for trees $T$ established by Cardoso, Jacobs, and Trevisan cardoso2017laplacian is tight.
Let $T_{n}$ be the tree constructed from the path $P_{3n}=v_{1}v_{2}\dots v_{3n}$, $n\geqslant 2$, by adding a leaf vertex adjacent to every vertex $v_{3i-1}$ for $2\leqslant i\leqslant n$ (See Figure 4). Root $T_{n}$ at $v_{3n}$ and run Algorithm 1 with $\alpha=-2$. The number of Laplacian eigenvalues in the interval $[2,n]$, that is $\nu(T)$, is equal to the number of non-negative vertices at the end of this run. One can verify that all the leaves and the vertices $v_{3i}$, $1\leqslant i\leqslant n$ end with negative values and the remaining vertices end with positive values. Hence $\nu(T_{n})=2n-1$. It is easy to check that $\gamma(T_{n})=n$.
$v_{1}$-$v_{2}1$$v_{3}$-$v_{4}1$$v_{5}$$v_{6}-1$$v_{7}$$v_{8}1$$v_{9}$-21-11-1
Figure 4: $T_{3}$ with inertia of vertices when rooted at $v_{9}$.
Therefore,
$\frac{\nu(T_{n})}{\gamma(T_{n})}=2-\frac{1}{n}.$
As $n$ increases, this ratio approaches 2.
7 Concluding remarks
In this paper we studied the relationship between the Laplacian spectrum and domination number in trees. The signless Laplacian matrix of a graph $G$ is $Q(G)=D(G)+A(G)$ and its spectrum is called the signless Laplacian spectrum of $G$. It is known that brouwer2011spectra a graph $G$ is bipartite if and only if the Laplacian spectrum and the signless Laplacian spectrum of $G$ are equal. Therefore our results on trees can also be stated in terms of their signless Laplacian spectrum.
Cardoso, Jacobs and Trevisan cardoso2017laplacian also proposed characterizing all graphs $G$ with $\mu(G)=\gamma(G)$. Characterizing trees $T$ with $\mu(T)=\gamma(T)$ is itself interesting and remains open.